# Trade-off of Security, Latency, and Throughput of the Nakamoto Consensus

### v3

@@ Appendix I Proof of Theorem 8
I.1. An Enlarged Violation Event
@@
Suppose $r^{\prime}\geq r\geq s\geq 0$. If an $r$-credible chain and an $r^{\prime}$-credible chain disagree on a target transaction that is public at time $s$, then
(226) $L_{s}+A_{s,r^{\prime}}\geq\hat{J}_{s,r}$
@@
If $s\geq r-\Delta$, we have $\hat{J}_{s,r}=0$, so (226) holds trivially. We assume $s<r-\Delta$ in the remainder of this proof. Let block $b$ be the first H-block mined after time $s$. Since the target transaction is public at time $s$, Definition 10 implies that there must be disagreements on all heights of jumpers mined from $s$ to $r-\Delta$, i.e., on no fewer than $J_{s,r-\Delta}$ heights. Let us denote the highest honest block before time $s$ as height $\eta$. Every unbalanced jumper mined during $(s,r-\Delta]$ must meet either one of the following two conditions:
@@
The adversary needs to mine at least on $\hat{J}_{s,r}$ heights of unbalanced jumpers. Due to the pre-mining phase, at time $s$, the highest A-block has a height $\eta+L_{s}$, which is $L_{s}$ higher than the highest honest block, which means at most $L_{s}$ of A-blocks can be mined before time $s$ on the heights of unbalanced jumpers. The rest A-blocks can only be mined after time $s$, the number of which is $A_{s,r^{\prime}}$. Therefore, the disagreement about the target transaction implies the inequality (226). Hence the lemma is proved. ∎
@@
(227) $F_{s,t}^{\epsilon}=\bigcup_{d\geq t+s}\left\{L_{s}+A_{s,d}\geq J_{s,d-\epsilon-\Delta}-X\right\}$
all $r^{\prime}$-credible chains agree on the target transaction for all $r^{\prime}\geq t+s$.
@@
(228) $L_{s}+A_{s,r^{\prime}}$ $\geq\hat{J}_{s,r}$
(229) $\geq\hat{J}_{s,r^{\prime}-\epsilon}$
(230) $\geq J_{s,r^{\prime}-\epsilon-\Delta}-X$
@@
Now suppose this lemma holds for $r^{\prime}\in[r,r+n\epsilon]$ for some positive integer $n$. We show the lemma also holds for $r^{\prime}\in[r,r+(n+1)\epsilon]$ as the following: First, let $r^{\prime\prime}=r+n\epsilon$. A repetition of the contradiction case above with $r$ replaced by $r^{\prime\prime}$ implies that the lemma also holds for $r^{\prime}\in[r^{\prime\prime},r^{\prime\prime}+\epsilon]$. Therefore lemma holds for $r^{\prime}\in[r,r+(n+1)\epsilon]$. The lemma is then established by induction on $n$.
@@
(231) $J_{s,d-\epsilon-\Delta}-X$ $\geq P_{s+\Delta,d-\epsilon-\Delta}-X.$
@@
(232) $F_{s,t}^{\epsilon}\subset\bigcup_{d\geq s+t}\left\{L_{s}+A_{s,d}\geq P_{s+\Delta,d-\epsilon-\Delta}-X\right\}.$
@@
(233) $P(F_{s,t}^{\epsilon})$ $\leq P(\bigcup_{d\geq s+t}\left\{L_{s}+A_{s,d}\geq P_{s+\Delta,d-\epsilon-\Delta}-X\right\})$
(234) $=P(\sup_{d\geq s+t}\{L_{s}+A_{s,d}-P_{s+\Delta,d-\epsilon-\Delta}\}+X\geq 0),$
@@
(235) $Y_{t}=\sup_{d\geq s+t}\{L_{s}+A_{s,d}-P_{s+\Delta,d-\epsilon-\Delta}\}.$
@@ I.2. Probability Bound
(238) $P(Y_{t}\leq-n)=\mathbb{E}[P(Y_{t}\leq-n|T)]$
(239) $=\mathbb{E}[P(\sup_{d\geq s+t}\{L_{s}+A_{s,d}-P_{T,d-\epsilon-\Delta}\}\leq-n|T)]$
(240) $=\mathbb{E}[P(\sup_{d\geq s+t}\{L_{s}+A_{s,d}-P^{\prime}_{0,d-\epsilon-\Delta-T}\}\leq-n|T)].$
@@
(241) $\geq\mathbb{E}[P(\sup_{d\geq s+t}\{L_{s}+A_{s,d}-P^{\prime}_{0,d-s-\epsilon-2\Delta}\}\leq-n|T)]$
(242) $=P(\sup_{d\geq s+t}\{L_{s}+A_{s,d}-P^{\prime}_{0,d-s-\epsilon-2\Delta}\}\leq-n)$
@@
(243) $\geq$ $P(\sup_{q\geq t}\{L_{s}+A^{\prime}_{0,q}-P^{\prime}_{0,q-\epsilon-2\Delta}\}\leq-n)$
(244) $=$ $P(\sup_{q\geq t}\{L_{s}+A^{\prime}_{0,t}+A^{\prime}_{t,q}-P^{\prime}_{0,t-\epsilon-2\Delta}-P^{\prime}_{t-\epsilon-2\Delta,q-\epsilon-2\Delta}\}\leq-n).$
We further replace $A^{\prime}_{t,q}$ by $A^{\prime\prime}_{0,q-t}$ in (244). We also note that $P^{\prime}_{t-2\Delta-\epsilon,q-2\Delta-\epsilon}$ is statistically dominated by $P^{\prime\prime}_{0,q-t}$. So, we rewrite (244) as
@@
(245) $\geq P\left(\sup_{q\geq t}\{L_{s}+A^{\prime}_{0,t}+A^{\prime\prime}_{0,q-t}-P^{\prime}_{0,t-\epsilon-2\Delta}-P^{\prime\prime}_{0,q-t}\}\leq-n\right)$
(246) $=P\left(L_{s}+A^{\prime}_{0,t}-P^{\prime}_{0,t-\epsilon-2\Delta}+\sup_{q\geq t}\{A^{\prime\prime}_{0,q-t}-P^{\prime\prime}_{0,q-t}\}\leq-n\right)$
(247) $=P\left(L_{s}+A^{\prime}_{0,t}-P^{\prime}_{0,t-\epsilon-2\Delta}+M\leq-n\right)$
(248) $=1-P(P_{0,t-\epsilon-2\Delta}^{\prime}\leq L_{s}+A_{0,t}^{\prime}+M+n-1)$
@@
(249) $M=\sup_{r\geq 0}\{A^{\prime\prime}_{0,r}-P^{\prime\prime}_{0,r}\}.$
As the maximum difference between two renewal processes, $M$ has the pmf (11) by (Kroese_difference, , Theorem 3.4) (see also (li2021close, , Sec. 6)). given by (11).
@@
$P(F_{s,t}^{\epsilon})$ $\leq\inf_{n\in\{1,2,...\}}\{P^{0}(X\geq n)$
(250) $+P(P_{0,t-\epsilon-2\Delta}^{\prime}\leq A_{0,t}^{\prime}+M+L_{s}+n-1)\}.$
@@
$\lim_{\epsilon\rightarrow 0}P(F_{s,t}^{\epsilon})$ $\leq\inf_{n\in\{1,2,...\}}\{P^{0}(X\geq n)$
(251) $+P(P_{0,t-2\Delta}^{\prime}\leq A_{0,t}^{\prime}+M+L_{s}+n-1)\}.$
@@
$\lim_{\epsilon\to 0}P(F_{s,t}^{\epsilon})\leq$ $\inf_{n\in\{1,2,...\}}\{P^{0}(X\geq n)$
(252) $+P(P_{0,t-2\Delta}^{\prime}\leq A_{0,t+2\Delta}^{\prime}+M+L+n-1)\}$
@@
(253) $P(P^{\prime}_{0,t}\leq i-1)$ $=P(i\Delta+X_{1}+\cdots+X_{i}>t)$
@@
$=$ $\sum_{i=0}^{\infty}\sum_{j=0}^{\infty}\sum_{k=0}^{\infty}P(A_{0,t+2\Delta}^{\prime}=i)P(M=j)P(L=k)$
(255) $\cdot P(P_{0,t-2\Delta}^{\prime}\leq i+j+k+n-1)$
$=$ $\sum_{i=0}^{\infty}\sum_{j=0}^{\infty}\sum_{k=0}^{\infty}f_{1}(i;a(t+2\Delta))e(j)e(k)$
(256) $\cdot(1-F_{2}(t-(i+j+k+n+2)\Delta;i+j+k+n,h)),$
where $f_{1}(\cdot;\omega)$ denotes the pmf of a Poisson distribution with mean $\omega$. (The cdf of $P^{\prime}_{0,t}$ has been evaluated in (li2021close, , Sec. 6.1).) It is easy to see that (256) is equal to $G(n,t)$ defined in (25).
@@
$\leq 1-P(A_{0,t+2\Delta}^{\prime}\leq i_{0})P(M\leq j_{0})P(L\leq k_{0})$
$\quad+P(P_{0,t-2\Delta}^{\prime}\leq A_{0,t+2\Delta}^{\prime}+M+L+n,A_{0,t+2\Delta}^{\prime}\leq i_{0},$
(258) $\quad M\leq j_{0},L\leq k_{0})$
(259) $\leq\overline{G}(n,t),$
@@
$\inf_{n\in\{1,2,\dots\}}\{P^{0}(X\geq n)+\overline{G}(n,t)\}$
(260) $=\inf_{n\in\{1,2,\dots\}}\{1-F(n-1)+\overline{G}(n,t)\}.$
@@
In this section, we lower bound the probability of safety violation achieved by a specific attack. The attack is the same as described in Appendix H, except that all H-blocks mined after $s-2\Delta$ are maximally delayed by $\Delta$, instead of after $s-\Delta$. Accordingly, the first new pacer is the first H-block mined after $s-2\Delta$.
Hence the probability that the adversary wins is lower bounded by
(261) $P$ $\left(\bigcup_{d\geq s+t}\left\{L_{s-2\Delta}+A_{s-2\Delta,d}\geq P^{\prime}_{d-s+\Delta}\geq 1\right\}\right).$
@@
$P\left(\bigcup_{d\geq s+t}\left\{L_{s-2\Delta}+A_{s-2\Delta,d}\geq P^{\prime}_{d-s+\Delta}\right\}\right)$
(262) $-P\left(\bigcup_{d\geq s+t}\left\{P^{\prime}_{d-s+\Delta}=0\right\}\right).$
The second term in (J) is easily determined as
(263) $P(\bigcup_{d\geq s+t}\left\{P^{\prime}_{d-s+\Delta}=0\right\})$ $=P(P^{\prime}_{t+\Delta}=0)$
(264) $=e^{-h(t+2\Delta)}.$
It remains to bound the first term in (J). Since $L_{s-2\Delta}$, $A_{s-2\Delta,d}$ and $P^{\prime}_{d-s+\Delta}$ are independent, we can replace $A_{s-2\Delta,d}$ with identically distributed and independent $A^{\prime}_{0,d-s+2\Delta}$ and replace $P^{\prime}_{d-s+\Delta}$ with $P^{\prime\prime}_{0,d-s+2\Delta}$ in (261) to obtain
$P$ $(\bigcup_{d\geq s+t}\left\{L_{s-2\Delta}+A_{s-2\Delta,d}\geq P^{\prime}_{d-s+\Delta}\right\})$
(265) $=P(\bigcup_{d\geq s+t}\left\{L_{s-2\Delta}+A^{\prime}_{0,d-s+2\Delta}\geq P^{\prime\prime}_{0,d-s+2\Delta}\right\})$
(266) $=P(\bigcup_{q\geq t}\left\{L_{s-2\Delta}+A^{\prime}_{0,t+\Delta}+A^{\prime}_{t+\Delta,q+2\Delta}\geq P^{\prime\prime}_{0,t+2\Delta}+P^{\prime\prime}_{t+2\Delta,q+2\Delta}\right\}).$
Evidently, $A^{\prime}_{t+\Delta,q+2\Delta}$ is identically distributed as $A^{\prime\prime}_{0,q-t+\Delta}$. Moreover, as a delayed renewal processes with identical inter-arrival times, $P^{\prime\prime}_{t+2\Delta,q+2\Delta}$ has statistically a larger delay than $P^{\prime}_{\Delta,q-t+\Delta}$. Thus $P^{\prime\prime}_{t+2\Delta,q+2\Delta}$ is stochastically dominated by $P^{\prime\prime}_{\Delta,q-t+\Delta}=P^{\prime\prime}_{0,q-t+\Delta}$. Hence (266) implies that
$P$ $(\bigcup_{d\geq s+t}\left\{L_{s-2\Delta}+A_{s-2\Delta,d}\geq P^{\prime}_{d-s+\Delta}\right\})$
(267) $\geq P(\bigcup_{q\geq t}\left\{L_{s-2\Delta}+A^{\prime}_{0,t+\Delta}+A^{\prime\prime}_{0,q-t+\Delta}\geq P^{\prime}_{0,t+2\Delta}+P^{\prime\prime}_{0,q-t+\Delta}\right\})$
(268) $\geq P(P^{\prime}_{0,t+2\Delta}\leq A^{\prime}_{0,t+2\Delta}+M+L_{s-2\Delta})$
where $M$ is defined in (249). The pmf of $M$ equals (11). Similar to (li2021close, ), $P(P^{\prime}_{0,t+2\Delta}\leq A^{\prime}_{0,t+2\Delta}+M+L_{s-2\Delta})$ can be calculated using the pmfs of the relevant independent random variables. Thus, given a latency $t$, the safety of a given transaction can be violated with a probability no less than given in (27).

### v5

@@ Appendix I Proof of Theorem 8
I-A An Enlarged Violation Event
@@
Suppose $r^{\prime}\geq r\geq s\geq 0$. If an $r$-credible chain and an $r^{\prime}$-credible chain disagree on a target block that is mined by time $s$, then
$L_{s}+A_{s,r^{\prime}}\geq\hat{J}_{s,r}$ (229)
@@
If $s\geq r-\Delta$, we have $\hat{J}_{s,r}=0$, so (229) holds trivially. We assume $s<r-\Delta$ in the remainder of this proof. Since the target block is mined by time $s$, there must be disagreements on all heights of jumpers mined from $s$ to $r-\Delta$, i.e., on no fewer than $J_{s,r-\Delta}$ heights. Let us denote the public height at time $s$ as height $\eta$. Every unbalanced jumper mined during $(s,r-\Delta]$ must meet either one of the following two conditions:
@@
The adversary needs to mine at least on $\hat{J}_{s,r}$ heights of unbalanced jumpers. Due to the pre-mining phase, at time $s$, the highest A-block has a height $\eta+L_{s}$, which is at most $L_{s}$ higher than the highest honest block, which means at most $L_{s}$ of A-blocks can be mined before time $s$ on the heights of unbalanced jumpers. The rest A-blocks can only be mined after time $s$, the number of which is $A_{s,r^{\prime}}$. Therefore, the disagreement about the target transaction implies the inequality (229). Hence the lemma is proved. ∎
@@
$F_{s,t}^{\epsilon}=\bigcup_{d\geq t+s}\left\{L_{s}+A_{s,d}\geq J_{s,d-\epsilon-\Delta}-X\right\}$ (230)
all $r^{\prime}$-credible chains agree on the target block for all $r^{\prime}\geq t+s$.
@@
$L_{s}+A_{s,r^{\prime}}$ $\geq\hat{J}_{s,r}$ (231)
$\geq\hat{J}_{s,r^{\prime}-\epsilon}$ (232)
$\geq J_{s,r^{\prime}-\epsilon-\Delta}-X$ (233)
@@
Now suppose this lemma holds for $r^{\prime}\in[r,r+n\epsilon]$ for some positive integer $n$. We show the lemma also holds for $r^{\prime}\in[r,r+(n+1)\epsilon]$ as the following: First, let $r^{\prime\prime}=r+n\epsilon$. A repetition of the contradiction case above with $r$ replaced by $r^{\prime\prime}$ implies that the lemma also holds for $r^{\prime}\in[r^{\prime\prime},r^{\prime\prime}+\epsilon]$. Therefore lemma holds for $r^{\prime}\in[r,r+(n+1)\epsilon]$. The lemma is then established by induction on $n$. ∎
@@
$J_{s,d-\epsilon-\Delta}-X$ $\geq P_{s+\Delta,d-\epsilon-\Delta}-X.$ (234)
@@
$F_{s,t}^{\epsilon}\subset\bigcup_{d\geq s+t}\left\{L_{s}+A_{s,d}\geq P_{s+\Delta,d-\epsilon-\Delta}-X\right\}.$ (235)
@@
$P(F_{s,t}^{\epsilon})$ $\leq P(\bigcup_{d\geq s+t}\left\{L_{s}+A_{s,d}\geq P_{s+\Delta,d-\epsilon-\Delta}-X\right\})$ (236)
$=P(\sup_{d\geq s+t}\{L_{s}+A_{s,d}-P_{s+\Delta,d-\epsilon-\Delta}\}+X\geq 0),$ (237)
@@
$Y_{t}=\sup_{d\geq s+t}\{L_{s}+A_{s,d}-P_{s+\Delta,d-\epsilon-\Delta}\}.$ (238)
@@ I-B Probability Bound
$P(Y_{t}\leq-n)$
$=\mathbb{E}[P(Y_{t}\leq-n|T)]$ (241)
$=\mathbb{E}[P(\sup_{d\geq s+t}\{L_{s}+A_{s,d}-P_{T,d-\epsilon-\Delta}\}\leq-n|T)]$ (242)
$=\mathbb{E}[P(\sup_{d\geq s+t}\{L_{s}+A_{s,d}-P^{\prime}_{0,d-\epsilon-\Delta-T}\}\leq-n|T)].$ (243)
@@
$\geq\mathbb{E}[P(\sup_{d\geq s+t}\{L_{s}+A_{s,d}-P^{\prime}_{0,d-s-\epsilon-2\Delta}\}\leq-n|T)]$ (244)
$=P(\sup_{d\geq s+t}\{L_{s}+A_{s,d}-P^{\prime}_{0,d-s-\epsilon-2\Delta}\}\leq-n)$ (245)
@@
$\geq P(\sup_{q\geq t}\{L_{s}+A^{\prime}_{0,q}-P^{\prime}_{0,q-\epsilon-2\Delta}\}\leq-n)$ (246)
$=P(\sup_{q\geq t}\{L_{s}+A^{\prime}_{0,t}+A^{\prime}_{t,q}-P^{\prime}_{0,t-\epsilon-2\Delta}-P^{\prime}_{t-\epsilon-2\Delta,q-\epsilon-2\Delta}\}\leq-n).$ (247)
We further replace $A^{\prime}_{t,q}$ by $A^{\prime\prime}_{0,q-t}$ in (247). We also note that $P^{\prime}_{t-2\Delta-\epsilon,q-2\Delta-\epsilon}$ statistically dominates $P^{\prime\prime}_{0,q-t}$. So, we rewrite (247) as
@@
$\geq P\left(\sup_{q\geq t}\{L_{s}+A^{\prime}_{0,t}+A^{\prime\prime}_{0,q-t}-P^{\prime}_{0,t-\epsilon-2\Delta}-P^{\prime\prime}_{0,q-t}\}\leq-n\right)$ (248)
$=P\left(L_{s}+A^{\prime}_{0,t}-P^{\prime}_{0,t-\epsilon-2\Delta}+\sup_{q\geq t}\{A^{\prime\prime}_{0,q-t}-P^{\prime\prime}_{0,q-t}\}\leq-n\right)$ (249)
$=P\left(L_{s}+A^{\prime}_{0,t}-P^{\prime}_{0,t-\epsilon-2\Delta}+M\leq-n\right)$ (250)
$=1-P(P_{0,t-\epsilon-2\Delta}^{\prime}\leq L_{s}+A_{0,t}^{\prime}+M+n-1)$ (251)
@@
$M=\sup_{r\geq 0}\{A^{\prime\prime}_{0,r}-P^{\prime\prime}_{0,r}\}.$ (252)
As the maximum difference between two renewal processes, $M$ has the pmf (12) by [22, Theorem 3.4] (see also [6, Sec. 6]). given by (12).
@@
$\begin{split}P(F_{s,t}^{\epsilon})&\leq\inf_{n\in\{1,2,...\}}\{P^{0}(X\geq n)+P(P_{0,t-\epsilon-2\Delta}^{\prime}\leq A_{0,t}^{\prime}+M+L_{s}+n-1)\}.\end{split}$ (253)
@@
$\begin{split}\lim_{\epsilon\rightarrow 0}P(F_{s,t}^{\epsilon})&\leq\inf_{n\in\{1,2,...\}}\{P^{0}(X\geq n)+P(P_{0,t-2\Delta}^{\prime}\leq A_{0,t}^{\prime}+M+L_{s}+n-1)\}.\end{split}$ (254)
@@
$\begin{split}\lim_{\epsilon\to 0}P(F_{s,t}^{\epsilon})\leq&\inf_{n\in\{1,2,...\}}\{P^{0}(X\geq n)+P(P_{0,t-2\Delta}^{\prime}\leq A_{0,t+2\Delta}^{\prime}+M+L+n-1)\}\end{split}$ (255)
@@
$P(P^{\prime}_{0,t}\leq i-1)$ $=P(i\Delta+X_{1}+\cdots+X_{i}>t)$ (256)
@@
$=$ $\sum_{i=0}^{\infty}\sum_{j=0}^{\infty}\sum_{k=0}^{\infty}P(A_{0,t+2\Delta}^{\prime}=i)P(M=j)P(L=k)\cdot P(P_{0,t-2\Delta}^{\prime}\leq i+j+k+n-1)$ (258)
$=$ $\sum_{i=0}^{\infty}\sum_{j=0}^{\infty}\sum_{k=0}^{\infty}f_{1}(i;a(t+2\Delta))e(j)e(k)\cdot(1-F_{2}(t-(i+j+k+n+2)\Delta;i+j+k+n,h)),$ (259)
where $f_{1}(\cdot;\omega)$ denotes the pmf of a Poisson distribution with mean $\omega$. (The cdf of $P^{\prime}_{0,t}$ has been evaluated in [6, Sec. 6.1].) It is easy to see that (259) is equal to $G(n,t)$ defined in (26).
@@
$\leq 1-P(A_{0,t+2\Delta}^{\prime}\leq i_{0})P(M\leq j_{0})P(L\leq k_{0})+P(P_{0,t-2\Delta}^{\prime}\leq A_{0,t+2\Delta}^{\prime}+M+L+n,A_{0,t+2\Delta}^{\prime}\leq i_{0},M\leq j_{0},L\leq k_{0})$ (261)
$\leq\overline{G}(n,t),$ (262)
@@
$\begin{split}&\inf_{n\in\{1,2,\dots\}}\{P^{0}(X\geq n)+\overline{G}(n,t)\}=\inf_{n\in\{1,2,\dots\}}\{1-F(n-1)+\overline{G}(n,t)\}.\end{split}$ (263)
@@
In this section, we lower bound the probability of safety violation achieved by a specific attack. The attack is the same as described in Appendix H. The first new pacer is the first H-block mined after $s$. Thus the probability that the adversary wins is lower bounded by
$P$ $\left(\bigcup_{d\geq s+t}\left\{L_{s}+A_{s,d}\geq P^{\prime}_{d-s-\Delta}\geq 1\right\}\right).$ (264)
@@
$\begin{split}P\left(\bigcup_{d\geq s+t}\left\{L_{s}+A_{s,d}\geq P^{\prime}_{d-s-\Delta}\right\}\right)-P\left(\bigcup_{d\geq s+t}\left\{P^{\prime}_{d-s-\Delta}=0\right\}\right).\end{split}$ (265)
The second term in (265) is easily determined as
$P(\bigcup_{d\geq s+t}\left\{P^{\prime}_{d-s-\Delta}=0\right\})$ $=P(P^{\prime}_{t-\Delta}=0)$ (266)
$=e^{-h(t-\Delta)}.$ (267)
It remains to bound the first term in (265). Since $L_{s}$, $A_{s,d}$ and $P^{\prime}_{d-s-\Delta}$ are independent, we can replace $A_{s,d}$ with identically distributed and independent $A^{\prime}_{0,d-s}$ and replace $P^{\prime}_{d-s-\Delta}$ with i.i.d. $P^{\prime\prime}_{0,d-s}$ in (264) to obtain
$P$ $(\bigcup_{d\geq s+t}\left\{L_{s}+A_{s,d}\geq P^{\prime}_{d-s-\Delta}\right\})$
$=P(\bigcup_{d\geq s+t}\left\{L_{s}+A^{\prime}_{0,d-s}\geq P^{\prime\prime}_{0,d-s}\right\})$ (268)
$=P(\bigcup_{q\geq t}\left\{L_{s}+A^{\prime}_{0,t}+A^{\prime}_{t,q}\geq P^{\prime\prime}_{0,t+\Delta}+P^{\prime\prime}_{t+\Delta,q}\right\}).$ (269)
Evidently, $A^{\prime}_{t,q}$ is identically distributed as $A^{\prime\prime}_{0,q-t}$. Moreover, as a delayed renewal processes with identical inter-arrival times, $P^{\prime\prime}_{t+\Delta,q}$ has statistically a larger delay than $P^{\prime}_{\Delta,q-t}$. Thus $P^{\prime\prime}_{t+\Delta,q}$ is stochastically dominated by $P^{\prime\prime}_{\Delta,q-t}=P^{\prime\prime}_{0,q-t}$. Hence (269) implies that
$P$ $(\bigcup_{d\geq s+t}\left\{L_{s}+A_{s,d}\geq P^{\prime}_{d-s-\Delta}\right\})$
$\geq P(\bigcup_{q\geq t}\left\{L_{s}+A^{\prime}_{0,t}+A^{\prime\prime}_{0,q-t}\geq P^{\prime}_{0,t+\Delta}+P^{\prime\prime}_{0,q-t}\right\})$ (270)
$\geq P(P^{\prime}_{0,t+\Delta}\leq A^{\prime}_{0,t}+M+L_{s})$ (271)
where $M$ is defined in (252). The pmf of $M$ equals (12). Similar to [6], $P(P^{\prime}_{0,t+\Delta}\leq A^{\prime}_{0,t}+M+L_{s})$ can be calculated using the pmfs of the relevant independent random variables. Thus, given a latency $t$, the safety of a given transaction can be violated with a probability no less than given in (28).